OpenStudy (anonymous):

Given: z=z(x,y) find... dz/dy partial if: x^4+3xy^2+5z=sin(z)

OpenStudy (anonymous):

can u explain?

OpenStudy (anonymous):

make it equal to 0 and differentiate with respect to y?

OpenStudy (anonymous):

just derivate implicitly. Think of x^4+3xy^2+5z=sin(z) like F(x,y,z) = x^4+3xy^2+5z-sin(z)=0 where z=z(x,y) sry, i errased the 1º answer, don't know what happened...

OpenStudy (anonymous):

so if z is a function of x and y how do i differentiate lets say -5z or -sin(z)?

OpenStudy (anonymous):

and no worries, i want to be able to get to answer myself

OpenStudy (anonymous):

-5z is -5z(x,y) so just use chain rule

OpenStudy (anonymous):

ohhhh ok i see, than this is easy, thanks

OpenStudy (anonymous):

wait distribute or treat them seperatly?

OpenStudy (anonymous):

still confused haha

OpenStudy (anonymous):

\[[-5z]'_{y} = -5*z'_{y}*dy\] chain rule...